Question 5 Tries remaining: 2 Points out of 10.00 Suppose you own an acre of land. You could grow crops on that land. The cost of seeds is $100. The crops you grow from those seeds will sell for $350. You could also rent the land to another farmer. The rent you could earn is $300. Calculate your economic profit. (Do not include a $ sign in your answer.) Flag question Answer: CheckQuestion 6 Tries remaining: 2 Points out of 10.00 Suppose the total cost function for a firm is given by: TC- 100+2q +0.5q2 Find the marginal cost function and then use that to determine the marginal cost of the 10th unit. (Do not include a $ sign in your response.) Flag question Answer Check0.5,.0.5 Suppose the production function for a firm is given by: q-4L K If the firm currently has 10 units of capital (K) Question Tries remaining: 2 and 10 units of labor (L), then calculate the Marginal Rate of Technical Substitution (MRTSLK) Points out of 10.00 (Round to the nearest two decimal places if necessary.) Flag question Answer: CheckQuestion 8 Tries remaining:2 and 10 units of labor (L), then calculate the Marginal Rate of Technical Substitution (MRTSLk). Points out of 10.00 (Round to the nearest two decimal places if necessary.) Suppose the production function for a firm is given by: q-2L +4K. If the firm currently has 20 units of capital (K) Flag question Answer: Check0.5,0.25 Question 9 Tries remaining: 1 Points out of 10.00 Suppose that a firm had a production function given by: q 2L KThe rental rate for the firm is $10 and the wage is $5. Solve the optimization condition for K and then fill in the value that appears in front of L. K; Flag question Round to the nearest 2 decimal places if necessary.) CheckO-L0k . The wage rate (w) s $ 10 and the rental rae Question 10 Suppose that a firm had a production function given by q=L 25K0 75 The wage rate w is $10 and the rental rate Tries remaining:2 r) is $5. Points out of 10.00 Calculate the amount of labor the firm would hire when it produces 400 units of output in a cost-minimizing way P Flag question (Round to the nearest 2 decimal places if necessary.) Answer: Check
Answer #1

1) Economic profit = Total revenue - Implicit cost -Explicit cost

Here Revenue = 350; Implict cost = 300; Explicit cost = 100

Economic profit = Total revenue - Implicit cost -Explicit cost

= 350 - 300 - 100

= -50

2)

TC = 100 + 2q + 0.5q2

MC = 100 *0 + 2 * 1 + 0.5*2q = 2 + q

Thus MC of 10th unit = 2 + 10 = 12